OpenStudy (anonymous):

Need help! Let F = Z3, and let W = { (a1, a2, a3, a4, a5) ∈ F5 | a1 + a5 = 0, a2 + 2a4 = 0 } (so that W is the set of all solutions to a homogeneous system of linear equations with coefficients from Z3 and is therefore a subspace of F5). Find a basis for W, and determine dim(W).

OpenStudy (anonymous):

basically, how do you do arithmetic in Z3? with {0,1,2}...supposedly you have to solve the system of equations but I dont know how to proceed

OpenStudy (anonymous):

\[a_1+a_5=0\] \[a_1=-a_5\] so choices here are \[a_1=a_5=0\] \[a_1=1,a_5=2\] or \[a_5=1, a_1=2\]

OpenStudy (anonymous):

ok, I understand that...to form a basis now, do we have to include all possible quintuples that offer solutions? because I did that and got more than 5 possibilities which cant be, needs to be dim5 or lower

OpenStudy (anonymous):

\[a_2+2a_4=0\] \[a_2=a_4=0\] or \[a_1=a_4=1\] or \[a_1=a_4=2\]

OpenStudy (anonymous):

you are working in \[\mathbb{Z_3}\] right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok now i have to think... first equation says \[a_1=-a_5\] second says \[a_2=a_4\]

OpenStudy (anonymous):

right, I took all possible scenarios and got 9 vectors, but it has to be 5 or less right? such as (0,0,0,0,0), (0,1,a3,1,0)...

OpenStudy (anonymous):

why does it have to be less than 5?

OpenStudy (anonymous):

you have a total of \[5^3=125\] possible vectors in this vector space.

OpenStudy (anonymous):

I thought the dimension has to be less than 5? because F^5

OpenStudy (anonymous):

i mean in the entire vector space. number of 5 tuples and three choices for each entry.

OpenStudy (anonymous):

Here is a reply I got about the question:Yes, the dimension of F^5 is 5 (basis e_1,e_2,e_3,e_4,e_5). This is true for any (every) field F. You just need to solve the system of linear equations ( the only difference is that the arithmetic is done in Z_3, not in R). The dimension of W is definitely less than that of F_5, since in a finite dimensional space, every proper subspace has dimension strictly less than the dimension of the whole space.

OpenStudy (anonymous):

I agree with you though, that's how I think

OpenStudy (anonymous):

oh yes the "dimension" is 5, but that does not mean there are five vectors only

OpenStudy (anonymous):

ok true perhaps, but there could only be 5 independent vectors right?

OpenStudy (anonymous):

we just have to come up with a basis, which should be easy if i could remember how to do this in general. the basis for the whole thing is the usual one \[\{(1,0,0,0,0),(0,1,0,0,0),(0,0,1,0,0),(0,0,0,1,0),(0,0,0,0,1)\}\] but we need a basis for this subspace

OpenStudy (anonymous):

right...we need to prove that basis is lin independent (i.e Ci=0 for CiVi) and that it generates W

OpenStudy (anonymous):

i have to think for a minute, but this should be straight forward. you have to come up with linearly independent vectors that span your whole subspace. there may only be one

OpenStudy (anonymous):

I found 9 vectors, (a3 can be anything right?) which would generate W, would the next step be to find if they are linearly independent?

OpenStudy (anonymous):

yes, and my guess is that there may be 3

OpenStudy (anonymous):

i mean 3 linearly independent ones

OpenStudy (anonymous):

how would you proceed with eliminating vectors?

OpenStudy (anonymous):

(0,0,1,0,0) (1,0,0,0,2) (0,1,0,1,0) is my guess

OpenStudy (anonymous):

ok how would you get (2,0,0,0,1)

OpenStudy (anonymous):

because as you said \[a_3\] can be anything so we have (0,0,1,0,0) as a basis

OpenStudy (anonymous):

i mean a basis vector

OpenStudy (anonymous):

\[2(1,0,0,0,2)=(2,0,0,0,1)\]

OpenStudy (anonymous):

ahhh right...ok

OpenStudy (anonymous):

So, find all possible solutions to W, then find the lin ind set of those vectors right?

OpenStudy (anonymous):

ok i think it is getting clearer to me, maybe. think of it this way: suppose you were working over the real numbers instead of Z3 then the equations would give you \[a_1=-a_5, a_2=-2a-4\] so you would have a basis that looked like \[(1,0,0,0,-1), (0,0,1,0,0), (0,2,0,1,0)\]

OpenStudy (anonymous):

typo there, i meant \[a_2=-2a_4\] of course

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and also \[(1,0,0,0,-1), (0,0,1,0,0), (0,2,0,-1,0)\]

OpenStudy (anonymous):

so it is a similar situation here but much easier because you only have three numbers to choose from . so i am going to bet that the basis is the one i wrote, and you can easily check that your nine vectors are linear combinations of the three

OpenStudy (anonymous):

i mean i would not bet a weeks pay on it, but i am pretty sure it is right

OpenStudy (anonymous):

and you could also show that the three are linearly independent. and that would not be much

OpenStudy (anonymous):

but it wont have -1 because that isn't an element of Z3

OpenStudy (anonymous):

right no -1 but that is ok because -1 = 2

OpenStudy (anonymous):

ok thanks a lot, I think I know how to proceedd...

OpenStudy (anonymous):

(0,0,1,0,0) (1,0,0,0,2) (0,1,0,1,0) becase a3 can be anything, a2 and a4 are equal, and whatever a1 is a5 is the other one

OpenStudy (anonymous):

yw